# How Far Are We From AGI?

### v1

@@ 4.6 Computing Platforms
The great power and efficiency of accelerators come with flexibility as well, which are granted by specially designed programming languages such as NVIDIA’s CUDA and AMD’s ROCm for more fine-grained control over thread utilization and computation logic. A bunch of works such as TVM (Chen et al., 2018) and MLC-LLM (MLC team, 2023) attempt to universalize the deployment of machine learning and deep learning models on everyone’s devices with compiler acceleration, which aims to maximize the potential of various accelerators. Research and engineering in AI hardware will likely drive the emergence of the next AI evolution, and we can expect that AGI systems need the next-generation hardware platforms that can break the current limitation and push the boundary of both computational and power efficiency to the next level.
@@ 4.7 The Future of AGI Systems
  • Systems co-designed with the model architecture to support efficient external resource acquisition. Being able to use diverse tools and acquire external knowledge is an indispensable requirement for future AGI systems. We can envision continual investment in developing and co-designing model-friendly tool interfaces (e.g. special APIs that differ from those used by humans and retrieval index that caters the model’s output patterns, etc) that can greatly improve the efficiency by which a model acquires external knowledge. One crucially desirable property of an AI system is life-long learning, which necessitates sophisticated memory and ability storage, a promising direction for system research.
@@
Moving towards the era of AGI, researchers and engineers can expect that investing in system research can enable even larger scale models with diversified data, a paradigm that has been shown in countless cases to be effective. Besides scaling, AGI system takes care of other aspects that are crucially important for practical deployment of these models such as privacy, trustworthiness, stability, and cost.
@@ 5 AGI Alignment: Ensuring AGI Meets Various Needs
The First Law: A robot may not injure a human being or, through inaction, allow a human being to come to harm. The Second Law: A robot must obey the orders given it by human beings except where such orders would conflict with the First Law. The Third Law: A robot must protect its own existence as long as such protection does not conflict with the First or Second Law.
@@
AGI alignment is a crucial technical approach for harnessing the capabilities of AGI, as discussed in the preceding sections, for practical applications in production and daily life. As shown in Figure 7, in this section, we begin by outlining expectations of AGI, addressing both its capabilities and the ethical considerations it entails. Subsequently, we explore current alignment techniques, which can be categorized into three distinct types: Online Human Supervision, Offline Human Supervision and Interactive Supervisions. Building on these discussions, we conclude by proposing a potential framework that classifies future AGI alignment strategies based on the type of AI interfaces summarized in Section § 3.
@@ 5.1 Expectations of AGI Alignment
To achieve this goal, researchers have proposed various approaches to AI alignment, such as value learning (Soares, 2016), inverse reinforcement learning (Hadfield-Menell et al., 2016), and cooperative inverse reinforcement learning (Hadfield-Menell et al., 2016). These methods aim to infer and align AI systems with human preferences and values. Additionally, it is crucial to develop ethical frameworks and guidelines that embrace a wide range of cultural, philosophical, and ethical perspectives, which will be discussed further in Section 5.3. This inclusivity helps mitigate biases and ensures a comprehensive representation of human values (Dignum, 2019).
@@
General Characteristics of AGI Alignments As discussed in sections 2, 3 and 4, AGI alignment should take AGI abilities into consideration, which includes AGI internal, AGI interface and AGI systems. Aligning AGI to human preferences also requires a thoughtful assessment of potential ethical issues at different scales. Prior research has attempted to outline and mitigate possible social risks of AI systems (Weidinger et al., 2021; Bender et al., 2021; Tamkin et al., 2021; Fjeld et al., 2020; Jobin et al., 2019). Building on previous works, we discuss several potential ethical issues that should be considered in AGI alignment.
@@
  • Privacy. AI systems can result in various types of digital privacy harms in the real world, arising from the unique capabilities of AI in emulating human- or superhuman-level performance at various tasks. According to previous work (Lee et al., 2024b; Das et al., 2023), on the one hand, AI systems can create new types of privacy risks. For example, they can yield risks including linking specific data points to an individual’s identity (Wiggers, 2021), combining various pieces of data about a person to make inferences beyond what is explicitly captured (Baraniuk, 2018), and inferring personality, social, and emotional attributes about an individual from their physical attributes (Levin, 2017). On the other hand, AI systems can also exacerbate many of the privacy risks that have existed even before the emergence of AI. For instance, AI systems can amplify secondary use risks by collecting user data for a different purpose without their consent (Long, 2021), and disclosure risks through sharing personal data to train models (Hodson, 2016), and intrusion risks via enabling centralized or ubiquitous surveillance infrastructures (Milmo, 2021). As AI starts to possess more human-level capabilities, such privacy risks will continue to exist in different forms.
@@ 5.2 Current Alignment Techniques
Aligning with Online Human Supervision Most current empirically verified LLMs alignment methods are in this group. These methods can generate online human supervision using techniques such as reinforcement learning or only inquiring about human supervision offline. We thus further divide these techniques in this group as with only human and offline human supervision. It is worth noting that methods in both subgroups have the potential to become a component of scalable oversight.
@@
The online supervision is acquired from the reward model during training. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) is one of the most prevalent methods for online supervised learning method. A variety of enhanced RLHF variants have also been proposed. The improving directions of RLHF are mainly from rule optimization, data utilization, and the theoretical justification aspect.
@@
  • Rule optimization. Sparrow (Glaese et al., 2022) incorporates adversarial probing and language-based rules into RLHF rewarding models. Bai et al. (2022a) investigate using pure RL to provide online human-level supervision for LLMs training and provide detailed explorations of the tradeoffs between output helpfulness and harmlessness.
@@
  • Optimal data utilization. Sensi (Liu et al., 2022c) tries to embed human value judgments into each step of language generation via a language model for both reward assignment (as the critic) and generation control (as the actor) during the generation. Baheti et al. (2023) focus on augmenting current training data by empowering different instances with varied weight for tuning models, taking the best advantage of the data w.r.t. their contribution to the language models.
@@
  • Theoretical justification. For improving current RLHF-related methods, f-DPG (Go et al., 2023) is framed as a generalization of RLHF to use any f-divergence to approximate any target distribution that can be evaluated, which differentiates it from previous method that can only fit KL divergence in the process. Zhu et al. (2023b) connect RLHF with max-entropy IRL (Ziebart et al., 2008) and propose a unified paradigm for such process with a sample complex bound for both situations.
@@
  • Text-based feedback signals. Text-based feedback signals involve converting human intents and preferences into text-based feedback to ensure alignment, extending the SFT process. These methods mainly expand from the improvement of training data. CoH (Liu et al., 2023f) is inspired by human learning processes, focusing on adjusting models based on successive outputs and summarized feedback from previous reasoning steps to fine-tune for predicting preferred outputs. RAFT (Dong et al., 2023b) uses a reward model to align model outputs with human preferences through SFT but in an offline manner. LIMA (Zhou et al., 2024a) aims to validate the assumption that LLMs acquire most knowledge during pre-training, requiring minimal instruction-tuning data to guide desirable output generation. ILF (Scheurer et al., 2023) introduces a three-stage process for modeling human preferences based on language feedback, showing parallels to Bayesian inference. Stable alignment (Liu et al., 2022c) learns alignment from multi-agent social interactions using a Sandbox simulator to optimize LLMs directly with preference data, avoiding reward hacking.
  • Ranking-based feedback signals. CRINGE (Adolphs et al., 2022) delves into negative examples that LLMs should steer clear of, while Xu et al. (2022) fine-tuning a model by training another model that generates toxic content. However, this methodology raises concerns regarding resource intensity and potential model quality and diversity degradation. Schick et al. (2021) put forth a methodology for identifying and generating text corresponding to toxic text types. SLiC (Zhao et al., 2023a) refines the probability of output sequences by aligning them with reference sequences using a variety of loss functions. RRHF (Yuan et al., 2023) generates supervision signals automatically for alignment through ranking results, whereas DPO (Rafailov et al., 2024) optimizes LLMs directly to align with human preferences, akin to RRHF but with a focus on maximizing reward and integrating KL divergence regularization. IPO (Azar et al., 2023) builds upon DPO by introducing a regularization term to stabilize the training process. Preference ranking optimization (PRO) (Song et al., 2023b) shares a similar approach with IPO and DPO in optimizing LLMs with ranking data but utilizes one positive and multiple negative samples rather than pairwise comparisons. Kahneman-Tversky Optimisation (KTO) (Ethayarajh et al., 2023) defines the loss function solely based on individual examples labeled as “good” or “bad” and does not necessitate pairwise preferences, making its training data more accessible.
@@
Aligning with Interactive Supervisions The ultimate goal for aligning models is regulating superhuman intelligence. A scalable aligning method is a promising mean that aims to address the challenge of overseeing complex tasks or superhuman models. By enabling relatively weak overseers, such as humans, to supervise complex tasks or systems using progressively evolved signals, scalable alignment offers a solution to tasks beyond human capabilities (Shen et al., 2023a).
@@
  • Through task decomposition. Various paradigms and strategies have been proposed to decompose complex tasks into simpler subtasks. Factored Cognition (Stiennon et al., 2020) involves breaking down a complex task into smaller, independent tasks processed simultaneously. Process Supervision (Lightman et al., 2023) fragments a task into sequential subtasks with supervision signals for each phase. Sandwiching (Bowman et al., 2022) delegates complex tasks to domain experts for resolution. IDA (Christiano et al., 2018) introduces an iterative distillation and amplification process that boosts the model’s capabilities through task decomposition. RRM (Leike et al., 2018) substitutes distilled imitation learning in IDA with reward modeling, optimizing the model using human-aligned signals and reinforcement learning. These methodologies aim to enhance collaboration between humans and agents for iterative improvement in solving complex tasks. SPIN (Chen et al., 2024a) decomposes the training objective into progressively evolved goals and leverages the iterative self-improvement to improve strong models with a weak one.
@@ 5.3 How to approach AGI Alignments
  • Interaction with tools and APIs. When interacting with tools and APIs, we mainly care about effectiveness, efficiency and some basic limiting rules in AGI alignment:
@@
Future AGI models are more capable at handling different tasks, and inevitably necessitate a significant increase in model parameters. To ensure their safe and effective deployment, we propose that research efforts focus on developing reliable, efficient, and transparent alignment techniques.
@@
  • Efficient alignments contribute to the blooming of AGI models. On the one hand, these methods rely heavily on the assumption that tasks can be parallelized (Segerie, 2023). This assumption may not always hold true, as some tasks, such as sorting algorithms, require sequential processing steps that cannot be fully decomposed into parallel parts, leading to extra processing time. On the other hand, the training stage is inevitable in these alignment methods. As the parameters scale becomes larger, this can be problematic when deploying alignment algorithms in real applications. Some recent works (Lin et al., 2023b) have started seeking solutions to reduce the overall training costs for aligning AI systems.
@@ 6.1 AGI Levels
This level of AGI usually performs better or on par with humans at specific benchmark tasks. Level-1 AGI represents the current state-of-the-art AI systems. For example, GPT-4 exhibits remarkable capabilities across many natural language tasks including language understanding, generating coherent and contextually relevant responses, often on par or superior to humans. These systems can often perform well given large enough human-curated datasets and are able to assist humans in certain domains. As indicated by the research (Bubeck et al., 2023), we are currently at this level of AGI in many domains.
@@
Given that we are still at the early stage of AGI, we acknowledge that our definitions might be high-level and abstract but nonetheless serve as theoretical guidelines. Therefore, to facilitate the understanding and better persuade the readers of the validity as well as generality of our definitions, we give several concrete examples in this section where we feature the main capabilities of each AI system as they evolve over the levels:
@@
      Self-driving (L2) cars are widely seen nowadays, facilitating not only drivers with a disability but also those who enjoy the semi-autonomous driving experience. In many closed facilities such as hotels, robots can reliably deliver food or items, which greatly preserves the privacy of the guests and saves the human cost. However, these semi-autonomous agents usually operate under a controlled environment or still require humans in case of emergency.
@@
      The whole city or even the globe is connected with the ultimate safety guarantees. A high-level planning constantly monitors all moving vehicles and can dynamically prioritize tasks that are of high importance such as emergency rescue and transportation coordination under special events. Personalized driving experience is also emphasized for those with different driving experience preferences. Finally, transportation is not limited to just cars but also flights and other robotics.
@@
Table 1: Comparison of AGI Levels and Their Characteristics. “L1”, “L2”, and “L3” refer to “Level 1”, “Level 2”, and “Level 3” of AGI respectively. For each of the main categories, we list several major conceptual criteria in terms of several categories that can be used to access whether we have reached a certain level of AGI.

### v2

@@ 4.6 Computing Platforms
The great power and efficiency of accelerators come with flexibility as well, which is granted by specially designed programming languages such as NVIDIA’s CUDA and AMD’s ROCm for more fine-grained control over thread utilization and computation logic. A bunch of works such as TVM (Chen et al., 2018) and MLC-LLM (MLC team, 2023) attempt to universalize the deployment of machine learning and deep learning models on everyone’s devices with compiler acceleration, which aims to maximize the potential of various accelerators. Research and engineering in AI hardware will likely drive the emergence of the next AI evolution, and we can expect that AGI systems need the next-generation hardware platforms that can break the current limitation and push the boundary of both computational and power efficiency to the next level.
@@ 4.7 The Future of AGI Systems
  • Systems co-designed with the model architecture to support efficient external resource acquisition. Being able to use diverse tools and acquire external knowledge is an indispensable requirement for future AGI systems. We can envision continual investment in developing and co-designing model-friendly tool interfaces (e.g. special APIs that differ from those used by humans and retrieval index that caters to the model’s output patterns, etc) that can greatly improve the efficiency by which a model acquires external knowledge. One crucially desirable property of an AI system is life-long learning, which necessitates sophisticated memory and ability storage, a promising direction for system research.
@@
Moving towards the era of AGI, researchers and engineers can expect that investing in system research can enable even larger-scale models with diversified data, a paradigm that has been shown in countless cases to be effective. Besides scaling, the AGI system takes care of other aspects that are crucially important for the practical deployment of these models such as privacy, trustworthiness, stability, and cost.
@@ 5 AGI Alignment: Ensuring AGI Meets Various Needs
The First Law: A robot may not injure a human being or, through inaction, allow a human being to come to harm. The Second Law: A robot must obey the orders given to it by human beings except where such orders would conflict with the First Law. The Third Law: A robot must protect its own existence as long as such protection does not conflict with the First or Second Law.
@@
AGI alignment is a crucial technical approach for harnessing the capabilities of AGI, as discussed in the preceding sections, for practical applications in production and daily life. As shown in Figure 7, in this section, we begin by outlining the expectations of AGI, addressing both its capabilities and the ethical considerations it entails. Subsequently, we explore current alignment techniques, which can be categorized into three distinct types: Online Human Supervision, Offline Human Supervision, and Interactive Supervision. Building on these discussions, we conclude by proposing a potential framework that classifies future AGI alignment strategies based on the type of AI interfaces summarized in Section § 3.
@@ 5.1 Expectations of AGI Alignment
To achieve this goal, researchers have proposed various approaches to AI alignment, such as value learning (Soares, 2016), inverse reinforcement learning (Hadfield-Menell et al., 2016), cooperative inverse reinforcement learning (Hadfield-Menell et al., 2016), and the most prevalent RLHF related strategies (Ouyang et al., 2022). These methods aim to infer and align AI systems with human preferences and values. Additionally, it is crucial to develop ethical frameworks and guidelines that embrace a wide range of cultural, philosophical, and ethical perspectives, which will be discussed further in Section 5.3. This inclusivity helps mitigate biases and ensures a comprehensive representation of human values (Dignum, 2019).
@@
General Characteristics of AGI Alignments As discussed in sections 2, 3, and 4, AGI alignment should take AGI abilities into consideration, which includes AGI internal, AGI interface, and AGI systems. Aligning AGI to human preferences also requires a thoughtful assessment of potential ethical issues at different scales. Prior research has attempted to outline and mitigate possible social risks of AI systems (Weidinger et al., 2021; Bender et al., 2021; Tamkin et al., 2021; Fjeld et al., 2020; Jobin et al., 2019). Building on previous works, we discuss several potential ethical issues that should be considered in AGI alignment.
@@
  • Privacy. AI systems can result in various types of digital privacy harms in the real world, arising from the unique capabilities of AI in emulating human- or superhuman-level performance at various tasks. According to previous work (Lee et al., 2024c; Das et al., 2023), on the one hand, AI systems can create new types of privacy risks. For example, they can yield risks including linking specific data points to an individual’s identity (Wiggers, 2021), combining various pieces of data about a person to make inferences beyond what is explicitly captured (Baraniuk, 2018), and inferring personality, social, and emotional attributes about an individual from their physical attributes (Levin, 2017). On the other hand, AI systems can also exacerbate many of the privacy risks that have existed even before the emergence of AI. For instance, AI systems can amplify secondary use risks by collecting user data for a different purpose without their consent (Long, 2021), and disclosure risks through sharing personal data to train models (Hodson, 2016), and intrusion risks via enabling centralized or ubiquitous surveillance infrastructures (Milmo, 2021). As AI starts to possess more human-level capabilities, such privacy risks will continue to exist in different forms.
@@ 5.2 Current Alignment Techniques
Aligning with Online Human Feedback Most current empirically verified LLMs alignment methods are in this group. These methods can help LLMs align with online human feedback using techniques such as reinforcement learning or only inquiring about human supervision offline (Tang et al., 2024a). We thus further divide these techniques in this group with only human and offline human supervision. It is worth noting that methods in both subgroups have the potential to become a component of scalable oversight.
@@
The online supervision is acquired from the reward model during training. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) is one of the most prevalent methods for online supervised learning method. A variety of enhanced RLHF variants have also been proposed. The improving directions of RLHF are mainly from reward modeling, optimization, data, and the self-improvement aspect.
@@
  • Reward Modeling. As the main supervision in the alignment process, reward modeling is a crucial way to improve the alignment techniques. Sparrow (Glaese et al., 2022) incorporates adversarial probing and language-based rules into RLHF rewarding models. Bai et al. (2022a) investigate using pure RL to provide online human-level supervision for LLMs training and provide detailed explorations of the tradeoffs between output helpfulness and harmlessness. Other techniques that unify both the reward and policy models have also emerged (Lee et al., 2024b), which broadens directions for aligning AI models. Another direction focusing on mitigating reward hacking or overoptimization issues, by updated evaluation protocols (Chen et al., 2024d), assembling multiple reward models (Ramé et al., 2024; Coste et al., 2023; Eisenstein et al., 2023), and refining the reward policy with synthetic data Pace et al. (2024).
  • Optimization. How to incorporate the supervision, either online or offline is an open question worth exploring. For example, Cheng et al. (2023) optimizes the reward model with the policy model simultaneously using min-max optimization. Recent works are exploring several alternatives for the RLHF method. DPO Rafailov et al. (2024) discards the reward model and optimizes for the final goal using the labeled preference in data. Muldrew et al. (2024) propose an improvement based on DPO with an active learning strategy. NashLLM (Munos et al., 2023) employs pairwise human feedback to train the policy model using Nash learning. ReMax (Li et al., 2023k) removes the value model in conventional reinforcement algorithms and introduces a novel variance reduction technique for stabilizing the optimization.
@@
  • Data. Sensi (Liu et al., 2022c) tries to embed human value judgments into each step of language generation via a language model for both reward assignment (as the critic) and generation control (as the actor) during the generation. Baheti et al. (2023) focus on augmenting current training data by empowering different instances with varied weight for tuning models, taking the best advantage of the data w.r.t. their contribution to the language models. To ensure continuous, high-quality data, AI-generated instances are utilized for adapting to RLHF (Lee et al., 2023) and, more recently, to DPO (Guo et al., 2024b)
@@
  • Self-Improvement. Strong AI models should learn how to improve themselves with or without external supervision. One of the most recent progress is the weak-to-strong generalization. For improving current RLHF-related methods, f-DPG (Go et al., 2023) is framed as a generalization of RLHF to use any f-divergence to approximate any target distribution that can be evaluated, which differentiates it from previous method that can only fit KL divergence in the process. Zhu et al. (2023b) connect RLHF with max-entropy IRL (Ziebart et al., 2008) and propose a unified paradigm for such a process with a sample complex bound for both situations.
@@
  • Text-based feedback signals. Text-based feedback signals involve converting human intents and preferences into text-based feedback to ensure alignment, extending the SFT process. These methods mainly expand from the improvement of training data. CoH (Liu et al., 2023f) is inspired by human learning processes, focusing on adjusting models based on successive outputs and summarized feedback from previous reasoning steps to fine-tune for predicting preferred outputs. RAFT (Dong et al., 2023c) uses a reward model to align model outputs with human preferences through SFT but in an offline manner. LIMA (Zhou et al., 2024a) aims to validate the assumption that LLMs acquire most knowledge during pre-training, requiring minimal instruction-tuning data to guide desirable output generation. ILF (Scheurer et al., 2023) introduces a three-stage process for modeling human preferences based on language feedback, showing parallels to Bayesian inference. Stable alignment (Liu et al., 2022c) learns alignment from multi-agent social interactions using a Sandbox simulator to optimize LLMs directly with preference data, avoiding reward hacking. SteerLM empowers end-users to control responses during inference by conditioning responses to conform to an explicitly defined multi-dimensional set of attributes (Dong et al., 2023b). CLP learns steerable models that effectively trade-off conflicting objectives at inference time based on techniques from multi-task training and parameter-efficient finetuning (Wang et al., 2024a).
[…]
@@
Scalable Oversight. The ultimate goal for aligning models is regulating superhuman intelligence. A scalable aligning method is a promising means that aims to address the challenge of overseeing complex tasks or superhuman models. By enabling relatively weak overseers, such as humans, to supervise complex tasks or systems using progressively evolved signals, scalable alignment offers a solution to tasks beyond human capabilities (Shen et al., 2023a).
@@
  • Through task decomposition. Various paradigms and strategies have been proposed to decompose complex tasks into simpler subtasks. Factored Cognition (Stiennon et al., 2020) involves breaking down a complex task into smaller, independent tasks processed simultaneously. Process Supervision (Lightman et al., 2023) fragments a task into sequential subtasks with supervision signals for each phase. Sandwiching (Bowman et al., 2022) delegates complex tasks to domain experts for resolution. IDA (Christiano et al., 2018) introduces an iterative distillation and amplification process that boosts the model’s capabilities through task decomposition. RRM (Leike et al., 2018) substitutes distilled imitation learning in IDA with reward modeling, optimizing the model using human-aligned signals and reinforcement learning. These methodologies aim to enhance collaboration between humans and agents for iterative improvement in solving complex tasks.
@@ 5.3 How to approach AGI Alignments
  • Interaction with tools and APIs. When interacting with tools and APIs, we mainly care about effectiveness, efficiency, and some basic limiting rules in AGI alignment:
@@
Future AGI models are more capable at handling different tasks and inevitably necessitate a significant increase in model parameters. To ensure their safe and effective deployment, we propose that research efforts focus on developing reliable, efficient, and transparent alignment techniques.
@@
  • Efficient alignments contribute to the blooming of AGI models. On the one hand, these methods rely heavily on the assumption that tasks can be parallelized (Segerie, 2023). This assumption may not always hold, as some tasks, such as sorting algorithms, require sequential processing steps that cannot be fully decomposed into parallel parts, leading to extra processing time. On the other hand, the training stage is inevitable in these alignment methods. As the parameters scale becomes larger, this can be problematic when deploying alignment algorithms in real applications. Some recent works (Lin et al., 2023b) have started seeking solutions to reduce the overall training costs for aligning AI systems.
@@ 6.1 AGI Levels
This level of AGI usually performs better or on par with humans at specific benchmark tasks (Bugaj and Goertzel, 2009). Level-1 AGI represents the current state-of-the-art AI systems. For example, GPT-4 exhibits remarkable capabilities across many natural language tasks including language understanding, generating coherent and contextually relevant responses, often on par or superior to humans. These systems can often perform well given large enough human-curated datasets and are able to assist humans in certain domains. As indicated by the research (Bubeck et al., 2023), we are currently at this level of AGI in many domains.
@@
Given that we are still at the early stage of AGI, we acknowledge that our definitions might be high-level and abstract but serve as theoretical guidelines. Therefore, to facilitate the understanding and better persuade the readers of the validity as well as generality of our definitions, we give several concrete examples in this section where we feature the main capabilities of each AI system as they evolve over the levels:
@@
      Self-driving (L2) cars are widely seen nowadays, facilitating not only drivers with disability but also those who enjoy the semi-autonomous driving experience. In many closed facilities such as hotels, robots can reliably deliver food or items, which greatly preserves the privacy of the guests and saves the human cost. However, these semi-autonomous agents usually operate under a controlled environment or still require humans in case of emergency.
@@
      The whole city or even the globe is connected with the ultimate safety guarantees. High-level planning constantly monitors all moving vehicles and can dynamically prioritize tasks that are of high importance such as emergency rescue and transportation coordination under special events. Personalized driving experience is also emphasized for those with different driving experience preferences. Finally, transportation is not limited to just cars but also flights and other robotics.
@@
Table 1: Comparison of AGI Levels and Their Characteristics. “L1”, “L2”, and “L3” refer to “Level 1”, “Level 2”, and “Level 3” of AGI respectively. For each of the main categories, we list several major conceptual criteria in terms of several categories that can be used to assess whether we have reached a certain level of AGI.